# Yajing Liu

### v1

@@ Abstract
We present ML-UCB, a generalized upper confidence bound algorithm that integrates arbitrary machine learning models into multi-armed bandit frameworks. A fundamental challenge in deploying sophisticated ML models for sequential decision-making is the lack of tractable concentration inequalities required for principled exploration. We overcome this limitation by directly modeling the learning curve behavior of the underlying estimator. Specifically, assuming the Mean Squared Error decreases as a power law in the number of training samples, we derive a generalized concentration inequality and prove that ML-UCB achieves sublinear regret. This framework enables the principled integration of any ML model whose learning curve can be empirically characterized, eliminating the need for model-specific theoretical analysis. We validate our approach through experiments on a collaborative filtering recommendation system using online matrix factorization with synthetic data designed to simulate a simplified two-tower model, demonstrating substantial improvements over LinUCB.
@@ 1 Introduction
We propose ML-UCB, a generalized UCB algorithm that integrates arbitrary machine learning models. Unlike previous methods, ML-UCB offers a model-agnostic framework that relies on the empirical learning curve, bridging the gap between theoretical rigor and the flexibility of arbitrary ML models. By leveraging the model’s learning curve, ML-UCB achieves faster regret convergence under specific conditions. Our contributions include: 1. A hybrid algorithm combining UCB with machine learning models. 2. A rigorous proof of sublinear regret for ML-UCB. 3. Experimental validation on the simulated dataset, demonstrating its effectiveness.
@@ 2 Algorithm
1. Model Training: Train the machine learning model on the available data. This step involves selecting a suitable model architecture and optimizing its parameters to minimize the training loss.
2. Learning Curve Estimation: Estimate the learning curve of the model by evaluating its Mean Squared Error (MSE) on a validation set. The learning curve provides insights into the model’s generalization ability as a function of the training sample size.
@@ 3. UCB Score Computation: Compute the UCB score for each arm using the generalized $\psi$-UCB formula:
Model Training Learning Curve Estimation UCB Score Computation Arm Selection Iterative Updates
@@
The flexibility of ML-UCB allows it to adapt to various machine learning models and datasets. For instance, in recommendation systems, collaborative filtering models can be used to predict user preferences.
@@ 4 Cumulant Generating Function and $\psi$-UCB
5 Calibrate Concentration Inequality for ML-model

### v2

@@ Abstract
We present ML-UCB, a generalized upper confidence bound algorithm that integrates arbitrary model-based estimators into multi-armed bandit frameworks. A fundamental challenge in deploying sophisticated ML models for sequential decision-making is the lack of tractable concentration inequalities required for principled exploration. We overcome this by directly modeling the learning curve behavior of the underlying estimator. Assuming the Mean Squared Error follows a power-law decay as training samples increase, we derive a generalized concentration inequality and prove ML-UCB achieves sublinear regret. This framework enables principled integration of any ML model whose learning curve can be empirically characterized, eliminating model-specific theoretical analysis. Our approach significantly reduces implementation complexity while saving compute and memory resources through its simple formula based on offline-trained parameters. Experiments on collaborative filtering with synthetic data demonstrate substantial improvements over LinUCB.
@@ 1 Introduction
We propose ML-UCB, a generalized UCB algorithm that integrates arbitrary machine learning models. Unlike previous methods, ML-UCB offers a model-agnostic framework that relies on the empirical learning curve, bridging the gap between theoretical rigor and the flexibility of arbitrary ML models. By leveraging the model’s learning curve, ML-UCB achieves faster regret convergence under specific conditions. Our contributions include: 1. A hybrid algorithm combining UCB with machine learning models. 2. A rigorous proof of sublinear regret for ML-UCB. 3. Rigorous validation on the simulated dataset, demonstrating its effectiveness.
@@ 2 Algorithm
1. Reward Model Validation: Validate the machine learning model’s performance on reward prediction using available data. This can be done either during the model training process or by testing a pre-trained model on a sample of arms. This approach treats the machine learning model as a black-box, abstracting away its internal micro-structure (e.g., the parameters of a neural network).
2. Learning Rate Estimation: Estimate the learning curve of the model by evaluating its Mean Squared Error (MSE) on a validation set. Importantly, the MSE here refers to the error of reward prediction, not the model’s training loss (e.g., cross-entropy for neural networks). The learning curve provides insights into the model’s generalization ability as a function of the training sample size. It is typically observed that the MSE decreases as a power law: the power is the learning rate $s$, which can be estimated by fitting a linear regression and used as a key parameter in the UCB score computation.
@@ 3. UCB Score Computation: Compute the UCB score for each arm using the generalized $\psi$-UCB formula:
Reward Model Validation Learning Rate Estimation UCB Score Computation Arm Selection Iterative Updates
@@
Compared with other approaches, this method offers several key advantages: it is more direct without requiring mathematical proofs based on estimations of model convergence speed from architectural details. It saves significant memory and computation since it relies on a simple formula based on an offline-trained parameter derived from the empirical learning curve. This parameter can be periodically tuned in response to changes in the learning curve, providing greater adaptability. ML-UCB adapts to various machine learning models and datasets, including deep neural networks, decision trees, and ensemble methods, making it versatile for diverse applications.
We call this algorithm ML-UCB, reflecting its core principle of adapting UCB through empirically fitted concentration inequalities.
@@ 4 Cumulant Generating Function and $\psi$-UCB
5 Calibrate Concentration Inequality for Model Based Estimators